OpenStudy (wyattp17):

Given the system of equations, what is the solution? x + 2y = 11 x - 2y = -1 {(-5, -3)} {(1, 1)} {(5, 3)}

OpenStudy (anonymous):

x + 2y = 11 x - 2y = -1 -----------add 2x + 0 = 10 2x = 10 x = 5 now sub 5 in for x in either of the original equations x + 2y = 11 5 + 2y = 11 2y = 11 - 5 2y = 6 y = 3 ANSWER : (5,3)